OpenStudy (anonymous):

Find the x and y intercepts of the line with the given equation... y=4x-1

OpenStudy (anonymous):

x int = 1/4 y int -1

OpenStudy (anonymous):

how did u get that answer?

OpenStudy (anonymous):

x int your imagine y=0 and for y int you put 0 for x

OpenStudy (anonymous):

x int = 0=4x-1 ==> x= 1/4 y int = y=4(0)-1 ==> y= -1